SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (22nd May, 2007)

Attendees: Debbie Bard, Fang Fang, Rahmat Rahmat.

Operations

  • Running has started again after the ROW and is going pretty smoothly.

GLT/online

  • Last Tuesday, while the magnet was ramped up and before beams were in the machine Fang installed the new GLT online libraries. Cosmics taken with them in looked fine and the data so far also looks fine so they've been left in place. See logbook entry.
  • The DCZ calibration/memory test software is still not working. Fang is debugging.

Data Quality and Monitoring

  • The TRG stripcharts are somehow different to the other DQ stripcharts and cause them to break when they're produced in Italy. Rahmat cannot reproduce the problem here, but it may be due to slightly different package versions or compilers. He is looking into it.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.